Give an example of a quadratic equation that satisfies each given condition:
1.) Has two equal real-number solutions.
2.) Has two distinct real-number solutions.
3.) Has non-real-number solutions.
Once again, help greatly appreciated. Thanks! :)

Answer by ad_alta(240)

1) x^2-2020202x+1020304030201=0
2) x^2+41500278x-117070685926648=0
3) 1010101x^2-x+2020202=0
You didn't think I would give you boring examples, did you?